The potential energy of a system of two atoms is given by the expression:

π‘ˆ = βˆ’ 𝐴 π‘Ÿ

2 ⁄ + 𝐡 π‘Ÿ

10 ⁄

A stable molecule is formed with release of 8.0 eV of energy when the interatomic distance isΒ 

2.8 Γ….Β 

i. Calculate A and B


1
Expert's answer
2021-01-14T10:37:31-0500

Answer

Given energy is

"E=8eV=12.8\times10^{-19}J"

Interatomic distance

"r=2.8A^°=2.8\times 10^{-10}m"

Potential energy is

"U=-Ar^2+Br^{10}"

Now force is given

"F=-\frac{dU}{dr}"

For equilibrium state

"F=0=-(-2Ar+10Br^9)"

By solving this

"r^8=\frac{A}{5B}"

"A=5Br^8"

"A=5\times (2.8\times10^{-10})^8B\\A=1.9\times10^{-76}B"

This relation shows value of A and B.